Suppose that people's heights (in centimeters) are normally distributed, with a mean of 170 and a standard deviation of 5. We find the heights of 80 people. (You may need to use the standard normal distribution table. Round your answers to the nearest whole number.)
(b) How many would you expect to be taller than 177 cm?

Answer by ewatrrr(24785)

Population: mean of 170 and a standard deviation of 5
Sample of 80
P(x>177) = P(z > 7/(5/sqrt(80)) = P(z > 7/.5590) = 1 - P(z ≤ 12.522) = basically 0%